OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? Answer 81 27 62 54

hero (hero):

Andres, re-check the previous question.

hero (hero):

x + y = 89 y = 2x + 8 y = 89 - x y = y 2x + 8 = 89 - x 3x = 97 x = 81/3 x = 27 y = 62

hero (hero):

x = girls y = boys